# Bridging Textual-Collaborative Gap through Semantic Codes for Sequential Recommendation

### v1

@@ 2.5. Discussion
To highlight the novelty and contributions of our approach, we compare CoCoRec with existing text-enhanced methods, as shown in Table 1. Firstly, previous methods typically adopt text representations and ID embeddings to capture textual and collaborative information, respectively. As a result, these two types of representations remain isolated and lack information sharing prior to their interaction (e.g., sum or attention mechanisms). In contrast, we establish a connection between them by adopting semantic codes to capture collaborative information. The connection between semantic codes and text representations allows codes to serve as bridges connecting the two types of information. Secondly, we adopt the code-guided cross attention mechanism for semantic fusion by considering semantic codes as the query to extract useful information from text representations, enabling flexible textual semantic integration. Thirdly, we introduce the masked code modeling task to capture the code relationships within individual items and improve the fusion between text representations and code embeddings.
@@
CoCoRec (ours) Text & Code Code-guided attention
@@ 3. Experiments
In this section, we empirically demonstrate the effectiveness of our proposed framework CoCoRec through extensive experiments and rigorous analysis.
@@ 3.1.1. Dataset
We conduct experiments on four subsets of the latest Amazon 2023 review data (Hou et al., 2024), i.e.,Musical Instruments”, “Video Games”, “Baby Products” and “Industrial Scientific”. To ensure a robust evaluation, we follow the same preprocessing steps (i.e., 5-core filtering) as described in prior studies (Zhou et al., 2020, 2022; Xu et al., 2024) to remove inactive users and unpopular items with less than five interactions. Next, we organize user behavior sequences in chronological order, limiting the maximum sequence length to 20 items. We categorize item attributes into five fields: title, brand, categories, features, and description, to capture fine-grained and rich semantic information. Each field’s content was truncated to a maximum of 512 tokens per item. The detailed statistics of the preprocessed datasets is presented in Table 2.
@@ 3.1.2. Baseline Models
Table 3. Performance comparisons among different methods. The best and second-best results are highlighted in bold and underlined font, respectively. “N@K” and “R@K” represent abbreviations of “NDCG@K” and “Recall@K”, respectively. “PQ” and “RQ” are short for “Product Quantization” and “Residual Quantization”. “Improv.” denotes the relative improvement ratios of CoCoRec over the top-performing baselines.
@@ 3.1.4. Implementation Details
For CoCoRec, we leverage Sentence-T5 (Ni et al., 2022) to encode the text attributes associated with each item as its text embeddings. Both the semantic-fusion module and the sequential recommender are configured with 2 layers and 2 attention heads, with a hidden dimension of 512. The embedding size is fixed at 128, while each attribute vector is quantized into 4 sub-vectors (i.e., $k=4$) with a codebook size of 256. The temperature coefficient $\tau$ and mask ratio $\rho$ are set to 0.07 and 0.5, respectively. We adopt the Adam optimizer for model training, with the learning rate tuned in {0.003, 0.001, 0.0005}. The dropout probability is tuned within {0.1, 0.2, 0.3, 0.4, 0.5} and the loss coefficients $\alpha$ and $\beta$ are searched in {0.2, 0.4, 0.6, 0.8} for optimal performance. To avoid overfitting, we utilize an early-stopping strategy, halting training when NDCG@10 shows no improvement over 10 consecutive epochs. We implement the majority of baseline models using the open-source recommendation system library RecBole (Zhao et al., 2021). For CL-based SR models, namely DuoRec and MAERec, we utilize the self-supervised learning framework SSLRec (Ren et al., 2024). For fair comparisons, we fix the embedding size of all models to 128 and perform the hyperparameter grid search for optimal results. Furthermore, all Transformer-based baseline models adopt the same recommendation backbone architecture as our CoCoRec for fairness.
@@ 3.2. Overall Performance
We compare CoCoRec with various baseline approaches on four public benchmarks and present the overall results in Table 3. From the results, we have the following findings:
@@
Finally, our proposed CoCoRec(PQ) achieves the best or second-best performance across all datasets, significantly outperforming the best baseline models in most cases. CoCoRec(RQ) also achieves competitive results but performs slightly worse than CoCoRec(PQ), which may be attributed to the fact that VQ-based semantic codes preserving semantics of different levels are more suitable for our framework than PQ-based codes implying semantics of different granularities. Different from existing text-enhanced methods, we propose leveraging semantic codes to bridge the gap between textual and collaborative information. With the specially designed semantic-fusion module and optimization objectives, including masked code modeling and masked sequence alignment, we achieve effective textual-collaborative semantic fusion for sequential recommendations, significant improving the model performance.
@@ 3.3. Ablation Study
To evaluate the contribution of each proposed component, we conduct ablation studies on the Instrument, Scientific, and Baby datasets. Table 4 presents the comparative results of the following four variants: (1) w/o $\mathcal{L}_{\text{MCM}}$ removes the masked code modeling (MCM) defined in Eq. (6). (2) w/o $\mathcal{L}_{\text{MSA}}$ without the masked sequence alignment (MSA) in Eq. (7). (3) w/o Text Emb substitutes text embeddings with semantic codes as cross-fusion layer input. (4) Random Code replaces the semantic codes derived from text embeddings with random codes.
The experimental results reveal that removing any component from CoCoRec consistently degrades model performance. This performance decline underscores the critical role of both proposed optimization objectives (i.e., MCM and MSA) in improving the representation learning of the overall framework. Furthermore, it demonstrates the significance of textual and collaborative information, as captured by text representations and code embeddings, respectively.
@@ 3.4.1. Impact of Item Attributes
To investigate how each item attribute affects the final performance, we conduct ablation studies on the Instrument and Scientific datasets by sequentially adding individual attributes. The results are reported in Table 5, from which we can find that text contents have a significant impact on the final performance. In general, richer attribute contents enable CoCoRec to extract more textual information beneficial for recommendation, allowing more refined item representations and user preference modeling.
@@ 3.4.2. Impact of Code Number
Table 6. Performance analysis of code number in CoCoRec. $k$, $m$ and $C$ denote the number of sub-vectors in vector quantization, the number of attributes and the codebook size of each sub-vector, respectively.
@@
We analyze the impact of code number on the final recommendation performance of CoCoRec by changing the number of sub-vectors in the vector quantization from 2 to 8. The results in Table 6 indicate that CoCoRec consistently benefits from increasing the code number. The reason is that a larger code number allows CoCoRec to capture textual semantics and model the item correlations more effectively in a broader representation space. However, increasing the number of codes also raises computational resource requirements. Therefore, it is essential to strike a balance between model performance and computational efficiency.
@@ 3.4.3. Impact of Code Type
To examine the impact of code types, we replace the original text-based semantic codes with three alternatives: random codes, codes derived from SASRec embeddings and mixed codes that combine equal proportions of text-based and SASRec-based codes. To ensure a fair comparison, the number of semantic codes in each case is kept consistent. As shown in Figure 2, the use of SASRec-based semantic codes leads to significant performance degradation. This is primarily because the collaborative information captured by SASRec-based codes is redundant, as CoCoRec is already capable of learning it. Our results indicate that models perform better when using more text semantic codes (Text ¿ Mix ¿ SASRec), underscoring the importance of textual semantic codes. Interestingly, random codes outperform SASRec-based codes, likely due to their stronger regularization effect.
3.4.4. Impact of Loss Coefficients
$\alpha$ and $\beta$ are key hyperparameters in CoCoRec, controlling the importance of the masked code modeling (MCM) and masked sequence alignment (MSA) objectives. We evaluate their impact by varying their values within {0.2, 0.4, 0.6, 0.8} on the Instrument and Scientific datasets. The results for Recall@10 and NDCG@10 are shown in Figure 3. On the Instrument dataset, CoCoRec’s performance initially improves but declines as $\alpha$ increases from 0.2 to 0.8, as moderate $\alpha$ values enhance semantic fusion while excessive values disrupt sequential pattern learning. On the Scientific dataset, the impact of $\alpha$ varies with $\beta$, indicating MCM’s influence is less predictable. For $\beta$, its effect depends on $\alpha$, with the optimal value shifting based on $\alpha$, highlighting the need to balance both hyperparameters for optimal performance.
@@ 3.4.5. Compatibility with ID-based Representations
In previous sections, we excluded unique item IDs and relied solely on text embeddings and their associated semantic codes to model items and user preferences. However, our proposed model, CoCoRec, is also compatible with ID-based representations. Below, we present a parameter-efficient approach for integrating unique item ID representations into our framework. To begin with, we pre-train CoCoRec using the original pipeline. Next, we introduce a raw item ID embedding table while freezing all other components of CoCoRec. The unique ID embeddings are then summed up with the semantic-fused item representations and fed as input to the sequential recommender. The model is then fine-tuned according to the recommendation objective $\mathcal{L}_{\text{CE}}$, as defined in Eq. (5), until convergence. This fine-tuning process is highly efficient since only the item embeddings are trainable. The experimental results are summarized in Table 7. The incorporation of unique item IDs further enhances the performance of CoCoRec. This improvement can be attributed to the incremental collaborative information provided by unique IDs, which enable CoCoRec to better distinguish between items. Unlike semantic codes, unique IDs offer a more flexible representation space, allowing for finer-grained differentiation.
@@
Table 7. Performance of CoCoRec empowered with ID-based representations on four datasets.
@@ 4.2. Text-Enhanced Recommendation
With the rapid development of various application platforms, a growing amount of item content information has become available. This has motivated researchers to explore its potential for improving the performance of sequential recommender systems. We refer to this approach as text-enhanced recommendation. Previous studies (Liu et al., 2021; Xie et al., 2022b; Xi et al., 2024; Gao et al., 2024; Xu et al., 2024; Liu et al., 2024a) have typically utilized text embeddings encoded by pre-trained language models (PLMs) to enhance item representations and improve behavior modeling. For instance, FDSA (Zhang et al., 2019) employs a dual-stream Transformer-based network to independently model item IDs and feature sequences. UniSRec (Hou et al., 2022) uses an MoE-enhanced adaptor to extract universal information from text embeddings through multi-domain pretraining. Another line of research (Petrov and Macdonald, 2024; Hu et al., 2024; Rajput et al., 2023) utilizes the discrete codes derived from the text embeddings in sequential or generative recommendation. For instance, VQRec (Hou et al., 2023) applies vector quantization to transform text embeddings into discrete item IDs, leveraging these quantized representations to model user preferences. MMSR (Hu et al., 2023) uses clustering to decompose original modality features into discrete semantic codes, treating these indices as nodes to build a dense item relation graph. Unlike the above methods, in CoCoRec, we leverage discrete semantic codes as bridges to achieve effective textual-collaborative semantic fusion.
@@ 5. Conclusion
In this work, we proposed CoCoRec to achieve effective textual and collaborative semantic fusion for sequential recommendation. The core idea of our approach is to bridge the textual-collaborative semantic gap through semantic codes. For this purpose, we construct multi-view text embeddings and corresponding semantic codes derived from vector quantization for fine-grained semantic utilization. Next, we design a code-guided semantic-fusion module to enable flexible textual information extraction and integration. Furthermore, we devise an optimization method via code masking, comprising masked code modeling and masked sequence alignment, to model code correlations within individual items and enhance sequence modeling. Extensive experiments and in-depth analysis on four latest benchmarks demonstrate the effectiveness of our proposed CoCoRec. In future work, we will leverage our method in other recommendation tasks (e.g., multi-modal recommendation and multi-domain recommendation).
Appendix A Time Complexity
Training. As outlined in Section 2, CoCoRec consists of two key components, the semantic-fusion module (SFM) and the sequential recommendation backbone (SR). For the SFM, the computational complexities of the self-attention layer, the cross-fusion layer, and the feed-forward network when processing a single item are $\mathcal{O}(M^{2}d)$, $\mathcal{O}(MKd)$ and $\mathcal{O}(Md^{2})$, respectively. Here, $d$ denotes the model dimension, while $M$ and $K$ represent the length of the semantic code sequence and item attribute sequence. Consequently, the overall time complexity for deriving a semantic-fused item representation is $\mathcal{O}(M^{2}d+MKd+Md^{2})$. For the SR, which utilizes a SASRec-like architecture, the time consumption of user behavior modeling is $\mathcal{O}(N^{2}d+Nd^{2})$, where $N$ is the length of the user interactions. Thus, the total training complexity of CoCoRec is $\mathcal{O}(N^{2}d+Nd^{2}+NM^{2}d+NMKd+NMd^{2})$. For brevity, the computation of training objectives is omitted. Notably, in practical scenarios, the code length $M$ and attribute number $K$ are typically much smaller than interaction sequence length $N$ (i.e., $M,K\ll N$). This ensures that the additional computational overhead introduced by CoCoRec remains acceptable for real-world applications.

### v2

@@ 2.5. Discussion
To highlight the novelty and contributions of our approach, we compare CCFRec with existing text-enhanced methods, as shown in Table 1. Firstly, previous methods typically adopt text representations and ID embeddings to capture textual and collaborative information, respectively. As a result, these two types of representations remain isolated and lack information sharing prior to their interaction (e.g., sum or attention mechanisms). In contrast, we establish a connection between them by adopting semantic codes to capture collaborative information. The connection between semantic codes and text representations allows codes to serve as bridges connecting the two types of information. Secondly, we adopt the code-guided cross attention mechanism for semantic fusion by considering semantic codes as queries to extract useful information from text representations, enabling flexible textual semantic integration. Thirdly, we introduce the masked code modeling task to capture the code relationships within individual items and improve the fusion between text representations and code embeddings.
@@
CCFRec (ours) Text & Code Code-guided attention
3. Time Complexity
Training. As outlined in Section 2, CCFRec consists of two key components, the semantic-fusion module (SFM) and the sequential recommendation backbone (SR). For the SFM, the computational complexities of the self-attention layer, the cross-fusion layer, and the feed-forward network when processing a single item are $\mathcal{O}(M^{2}d)$, $\mathcal{O}(MKd)$ and $\mathcal{O}(Md^{2})$, respectively. Here, $d$ denotes the model dimension, while $M$ and $K$ represent the length of the semantic code sequence and item attribute sequence. Consequently, the overall time complexity for deriving a semantic-fused item representation is $\mathcal{O}(M^{2}d+MKd+Md^{2})$. For the SR, which utilizes a SASRec-like architecture, the time consumption of user behavior modeling is $\mathcal{O}(N^{2}d+Nd^{2})$, where $N$ is the length of the user interactions. Thus, the total training complexity of CCFRec is $\mathcal{O}(N^{2}d+Nd^{2}+NM^{2}d+NMKd+NMd^{2})$. For brevity, the computation of training objectives is omitted. Notably, in practical scenarios, the code length $M$ and attribute number $K$ are typically much smaller than interaction sequence length $N$ (i.e., $M,K\ll N$). This ensures that the additional computational overhead introduced by CCFRec remains acceptable for real-world applications.
@@ 4. Experiments
In this section, we empirically demonstrate the effectiveness of our proposed framework CCFRec through extensive experiments and rigorous analysis.
@@ 4.1.1. Dataset
We conduct experiments on four subsets of the latest Amazon 2023 review data (Hou et al., 2024), i.e.,Musical Instruments”, “Video Games”, “Baby Products” and “Industrial Scientific”. To ensure a robust evaluation, we follow the same preprocessing steps (i.e., 5-core filtering) as described in prior studies (Zhou et al., 2020, 2022; Xu et al., 2024) to remove inactive users and unpopular items with less than five interactions. Next, we organize user behavior sequences in chronological order, limiting the maximum sequence length to 20 items. We categorize item attributes into five fields: title, brand, categories, features, and description, to capture fine-grained and rich semantic information. Each field’s content is truncated to a maximum of 512 tokens per item. The detailed statistics of the preprocessed datasets are presented in Table 2.
@@ 4.1.2. Baseline Models
Table 3. Performance comparisons among different methods. The best and second-best results are highlighted in bold and underlined font, respectively. “N@K” and “R@K” represent abbreviations of “NDCG@K” and “Recall@K”, respectively. “PQ” and “RQ” are short for “Product Quantization” and “Residual Quantization”. “Improv.” denotes the relative improvement ratios of CCFRec over the top-performing baselines.
@@ 4.1.4. Implementation Details
For CCFRec, we leverage Sentence-T5 (Ni et al., 2022) to encode the text attributes associated with each item as its text embeddings. Both the semantic-fusion module and the sequential recommender are configured with 2 layers and 2 attention heads, with a hidden dimension of 512. The embedding size is fixed at 128, while each attribute vector is quantized into 4 sub-vectors (i.e., $k=4$) with a codebook size of 256. The temperature coefficient $\tau$ and mask ratio $\rho$ are set to 0.07 and 0.5, respectively. We adopt the Adam optimizer for model training, with the learning rate tuned in {0.003, 0.001, 0.0005}. The dropout probability is tuned within {0.1, 0.2, 0.3, 0.4, 0.5} and the loss coefficients $\alpha$ and $\beta$ are searched in {0.2, 0.4, 0.6, 0.8} for optimal performance. To avoid overfitting, we utilize an early-stopping strategy, halting training when NDCG@10 in the validation set shows no improvement over 10 consecutive epochs. We implement the majority of baseline models using the open-source recommendation system library RecBole (Zhao et al., 2021, 2022; Xu et al., 2023). For CL-based SR models, namely DuoRec and MAERec, we utilize the self-supervised learning framework SSLRec (Ren et al., 2024). For fair comparisons, we fix the embedding size of all models to 128 and perform the hyperparameter grid search for optimal results. Furthermore, all Transformer-based baseline models adopt the same recommendation backbone architecture as our CCFRec for fairness.
@@ 4.2. Overall Performance
We compare CCFRec with various baseline approaches on four public benchmarks and present the overall results in Table 3. From the results, we have the following findings:
@@
Finally, our proposed CCFRec(PQ) achieves the best or second-best performance across all datasets, significantly outperforming the best baseline models in most cases. CCFRec(RQ) also achieves competitive results but performs slightly worse than CCFRec(PQ), which may be attributed to the fact that PQ-based semantic codes preserving semantics of different levels are more suitable for our framework than RQ-based codes implying semantics of different granularities. Different from existing text-enhanced methods, we propose leveraging semantic codes to bridge the gap between textual and collaborative information. With the specially designed semantic-fusion module and optimization objectives, including masked code modeling and masked sequence alignment, we achieve effective textual-collaborative semantic fusion for sequential recommendations, significantly improving the model performance.
@@ 4.3. Ablation Study
To evaluate the contribution of each proposed component, we conduct ablation studies on the Instrument, Scientific, and Baby datasets. Table 4 presents the comparative results of the following six variants: (1) w/o $\mathcal{L}_{\text{MCM}}$ removes the masked code modeling (MCM) defined in Eq. (6). (2) w/o $\mathcal{L}_{\text{MSA}}$ without the masked sequence alignment (MSA) in Eq. (7). (3) w/o Text Emb substitutes text embeddings with semantic codes as cross-fusion layer input. (4) Random Code replaces the semantic codes derived from text embeddings with random codes. (5) Global Emb encodes all text attributes into a single global embedding, replacing the individual attribute embeddings (6) w/o CA replaces the SFM with mean pooling of all associated text representations and code embeddings to generate the final item representation.
The experimental results reveal that removing any component from CCFRec consistently degrades model performance. This performance decline underscores the critical role of both proposed optimization objectives (i.e., MCM and MSA) in improving the representation learning of the overall framework. Furthermore, it demonstrates the significance of textual and collaborative information, as captured by text representations and code embeddings, respectively.
@@ 4.4.1. Impact of Item Attributes
To investigate how each item attribute affects the final performance, we conduct ablation studies on the Instrument and Scientific datasets by sequentially adding individual attributes. The results are reported in Table 5, from which we can find that text contents have a significant impact on the final performance. In general, richer attribute contents enable CCFRec to extract more textual information beneficial for recommendation, allowing more refined item representations and user preference modeling.
@@ 4.4.2. Impact of Code Number
Table 6. Performance analysis of code number in CCFRec. $k$, $m$ and $C$ denote the number of sub-vectors in vector quantization, the number of attributes and the codebook size of each sub-vector, respectively.
@@
We analyze the impact of code number on the final recommendation performance of CCFRec by changing the number of sub-vectors in the vector quantization from 2 to 8. The results in Table 6 indicate that CCFRec consistently benefits from increasing the code number. The reason is that a larger code number allows CCFRec to capture textual semantics and model the item correlations more effectively in a broader representation space. However, increasing the number of codes also raises computational resource requirements. Therefore, it is essential to strike a balance between model performance and computational efficiency.
@@ 4.4.3. Impact of Code Type
To examine the impact of code types, we replace the original text-based semantic codes with three alternatives: random codes, codes derived from SASRec embeddings and mixed codes that combine equal proportions of text-based and SASRec-based codes. To ensure a fair comparison, the number of semantic codes in each case is kept consistent. As shown in Figure 2, the use of SASRec-based semantic codes leads to significant performance degradation. This is primarily because the collaborative information captured by SASRec-based codes is redundant, as CCFRec is already capable of learning it. Our results indicate that models perform better when using more text semantic codes (Text ¿ Mix ¿ SASRec), underscoring the importance of textual semantic codes. Interestingly, random codes outperform SASRec-based codes, likely due to their stronger regularization effect.
4.4.4. Impact of Loss Coefficients
$\alpha$ and $\beta$ are key hyperparameters in CCFRec, controlling the importance of the masked code modeling (MCM) and masked sequence alignment (MSA) objectives. We evaluate their impact by varying their values within {0.2, 0.4, 0.6, 0.8} on the Instrument and Scientific datasets. The results for Recall@10 and NDCG@10 are shown in Figure 3. On the Instrument dataset, CCFRec’s performance initially improves but declines as $\alpha$ increases from 0.2 to 0.8, as moderate $\alpha$ values enhance semantic fusion while excessive values disrupt sequential pattern learning. On the Scientific dataset, the impact of $\alpha$ varies with $\beta$, indicating MCM’s influence is less predictable. For $\beta$, its effect depends on $\alpha$, with the optimal value shifting based on $\alpha$, highlighting the need to balance both hyperparameters for optimal performance.
@@ 4.4.5. Compatibility with ID-based Representations
In previous sections, we excluded unique item IDs and relied solely on text embeddings and their associated semantic codes to model items and user preferences. However, our proposed model, CCFRec, is also compatible with ID-based representations. Below, we present a parameter-efficient approach for integrating unique item ID representations into our framework. To begin with, we pre-train CCFRec using the original pipeline. Next, we introduce a raw item ID embedding table while freezing all other components of CCFRec. The unique ID embeddings are then summed up with the semantic-fused item representations and fed as input to the sequential recommender. The model is then fine-tuned according to the recommendation objective $\mathcal{L}_{\text{CE}}$, as defined in Eq. (5), until convergence. This fine-tuning process is highly efficient since only the item embeddings are trainable. The experimental results are summarized in Table 7. The incorporation of unique item IDs further enhances the performance of CCFRec. This improvement can be attributed to the incremental collaborative information provided by unique IDs, which enable CCFRec to better distinguish between items. Unlike semantic codes, unique IDs offer a more flexible representation space, allowing for finer-grained differentiation.
@@
4.4.6. Analysis of the balance between collaborative and textual signals.
To validate our model’s ability to bridge the textual and collaborative gap, we visualize its top-2000 similar items via t-SNE and use blue and red colors to represent the collaboratively similar and textually similar items, respectively. We present the examples of three items in Figure 4. The results reveal that CCFRec achieves a balanced distribution of textually and collaboratively similar items, confirming its ability to harmonize both information effectively.
Table 7. Performance of CCFRec empowered with ID-based representations on four datasets.
@@
Figure 4. Visualization of similar items with different signals on Instrument. The left, middle, and right panels display t-SNE visualizations of the top-2000 similar items for a specific item derived from the SASRec embeddings, text embeddings, and CCFRec’s fused representations, respectively.
@@ 5.2. Text-Enhanced Recommendation
With the rapid development of various application platforms, a growing amount of item content information has become available. This has motivated researchers to explore its potential for improving the performance of sequential recommender systems. We refer to this approach as text-enhanced recommendation. Previous studies (Liu et al., 2021; Xie et al., 2022b; Xi et al., 2024; Gao et al., 2024; Xu et al., 2024; Liu et al., 2024a) have typically utilized text embeddings encoded by pre-trained language models (PLMs) to enhance item representations and improve behavior modeling. For instance, FDSA (Zhang et al., 2019) employs a dual-stream Transformer-based network to independently model item IDs and feature sequences. UniSRec (Hou et al., 2022) uses an MoE-enhanced adaptor to extract universal information from text embeddings through multi-domain pretraining. Another line of research (Petrov and Macdonald, 2024; Hu et al., 2024; Rajput et al., 2023; Liu et al., 2024c) utilizes the discrete codes derived from the text embeddings in sequential or generative recommendation. For instance, VQRec (Hou et al., 2023) applies vector quantization to transform text embeddings into discrete item IDs, leveraging these quantized representations to model user preferences. MMSR (Hu et al., 2023) uses clustering to decompose original modality features into discrete semantic codes, treating these indices as nodes to build a dense item relation graph. Unlike the above methods, in CCFRec, we leverage discrete semantic codes as bridges to achieve effective textual-collaborative semantic fusion.
@@ 6. Conclusion
In this work, we proposed CCFRec to achieve effective textual and collaborative semantic fusion for sequential recommendation. The core idea of our approach is to bridge the textual-collaborative semantic gap through semantic codes. For this purpose, we construct multi-view text embeddings and corresponding semantic codes derived from vector quantization for fine-grained semantic utilization. Next, we design a code-guided semantic-fusion module to enable flexible textual information extraction and integration. Furthermore, we devise an optimization method via code masking, comprising masked code modeling and masked sequence alignment, to model code correlations within individual items and enhance sequence modeling. Extensive experiments and in-depth analysis on four latest benchmarks demonstrate the effectiveness of our proposed CCFRec. In future work, we will leverage our method in other recommendation tasks (e.g., multi-modal recommendation and multi-domain recommendation).
Acknowledgements.
This work was partially supported by National Natural Science Foundation of China under Grant No. 92470205 and 62222215, Beijing Municipal Science and Technology Project under Grant No. Z231100010323009, and Beijing Natural Science Foundation under Grant No. L233008. Xin Zhao is the corresponding author.